A boy purchased a party length sandwich 61 inches long. He wants to cut into three pieces so that the middle piece is 5 inches longer than the shortest piece and the shortest piece is 8 inches shorter than the longest piece. How long should the three pieces be?
Answer by Boreal(15235)

x=shortest piece
x+5=middle piece
x+8=longest piece
Add them and see them equal to 61 inches.
3x+13=61
3x=48; x=16 inches
the pieces are 16,21,and 24 inches, and they add to 61 inches.